# Efficient and Effective Adaptation of Multimodal Foundation Models in Sequential Recommendation

### v1

@@ III-E LayerDrop
Additionally, to further enhance network efficiency and address issues of layer redundancy, we introduce LayerDrop techniques, similar to those in [50, 57], to reduce the number of SAN blocks and align asymmetrical MFM. Specifically, for symmetrical MFM, we group two transformer blocks together and drop the first hidden state to the SANs, which can save half the number of SANBs. In the case of asymmetrical MFM, to address the imbalance between the layers of LLM and vision encoders, we apply a simple yet effective group LayerDrop technique that groups $k$ layers evenly from the hidden states of the LLM. Suppose there are $L^{\text{text}}$ and $L^{\text{image}}$ layers in the text encoder and image encoder, respectively. In scenarios of asymmetrical adaptation, the text encoder is an LLM, thus $L^{\text{text}}\gg L^{\text{image}}$. For image encoders, we continue using the same LayerDrop technique as with symmetrical structures, where we drop half the number of layers. In this paper, we design to select the largest $k$ to ensure the maximum range of cross-layer interaction within the entire LLM. The number $k$ can be represented as follows:
$\max_{k}\left(L^{\text{text}}-k\cdot\frac{L^{\text{image}}}{2}\geq 1\right)$ (5)
In Section VI-E, we further investigate various LayerDrop schemes, showcasing the performance impact of dropping different numbers of blocks in both symmetrical and asymmetrical structures.
@@ III-G Training Objectives
In terms of training details, we exploit the in-batch debiased Cross-Entropy loss function $\mathcal{L}_{CE}$ [60, 5, 61] widely adopted in both academic literature and industrial systems.
@@ IV Efficiency Analysis
To simplify the theoretical analysis, we consider a scenario involving large foundation models comprising $n$ transformer blocks, where the $i$-th transformer block is denoted by $TRM_{i}$. For each transformer block, all PEFT methods, including Adapter, LoRA, and IISAN, will place a trainable PEFT module denoted by $peft$. We assume that $TRM\gg peft$ according to the definition of parameter efficiency.
@@ V Experiment Setup
TABLE I: Dataset Description
@@ V-A Dataset
In this study, each item in a user sequence is represented by its raw modality information. To assess our methods on items that include both image and text modalities, we utilize the Amazon review datasets [62]. Specifically, we employ three widely used datasets from this collection: ”Industrial and Scientific,” ”Musical Instruments,” and ”Office Products.” These datasets were adopted following the preprocessing steps outlined in prior works [18, 5, 13, 63]. Furthermore, to explore the performance of IISAN-Versa compared with other recommendation approaches with a fair comparison, we also incorporate Microlens [24], a public multimodal video recommendation benchmark that includes raw text, images, and video content. Since Microlens is a dataset rich in modality information, we also explore multimodal text scenarios within this public benchmark.
@@ V-B Multimodal Text Generation
To explore the potential of IISAN-Versa in adapting multiple scenarios, such as multimodal text scenarios containing cover text and corresponding video texts, we utilize the pretrained LLaVA [^3] https://github.com/haotian-liu/LLaVA model to generate captions for image and video content from the public benchmark Microlens datasets, as illustrated in Figure 3. For this process, we input the cover image and extract three evenly spaced frames from each video. In terms of caption models, we use ”llava-hf/llava-v1.6-mistral-7b-hf” for cover image captioning. For video captioning, we apply ”liuhaotian/llava-v1.6-34b” due to the higher complexity of videos; a larger caption model provides a more accurate representation of the video content. Our video captioning strategies follow the preprocessing methods used in Video generation OpenSora [^4] https://github.com/hpcaitech/Open-Sora/. To enhance LLaVA’s ability to produce effective representations for postprocessing, we incorporate in-context learning by including concrete examples in the prompt. We adopt the prompt from the Sora [^5] https://openai.com/index/sora/ project for video prompts and use the prompt from DALLE2666https://openai.com/index/dall-e-2/ for cover image prompts.
@@
Based on previous research [64, 5, 13, 18], our methodology employs a leave-one-out evaluation strategy. In this approach, the final item in the interaction sequence is reserved for testing, the penultimate item is used for validation, and all preceding items are utilized for training. We evaluate our model using HR@10 (Hit Ratio) and NDCG@10 (Normalized Discounted Cumulative Gain), consistent with the evaluation methodologies in [64, 5, 13], as the key performance metrics. Unless stated otherwise, all results reported refer to the test set. Furthermore, it is crucial to mention that the predicted item is evaluated against the complete set of items [65].
TABLE II: Description of the pre-trained models used in this research. L denotes the number of Transformer blocks, H represents the hidden dimension of the transformer layer, B indicates the number of parameters in billions, and M indicates the number of parameters in millions.
@@
Text BERT-base L=12, H=768 110M
@@
V-D Implementation Details.
The pretrained models from the Huggingface platform [66], detailed in Table II, are employed as text and image encoders in this research. Note that due to the limited computing resources, we adopt the Llama-3-70b with GPTQ 4-bit quantization. We experiment with hidden representation dimensions for the sequential encoder from {32, 64, 128} and set it to 64. We fixed the sequential encoder to include 2 Transformer blocks and 2 attention heads following [18, 5]. Using the Adam optimizer without weight decay, we tested learning rates from 1e-5 to 1e-3 while maintaining a dropout probability of 0.1. We explored batch sizes from {8, 16, 32, 64, 512}, selecting the largest viable option to optimize GPU memory, and searched for SANB hidden dimensions from {8, 16, 32, 64, 128, 512, 1024, 2048}. Hyperparameters were searched separately for the Microlens and Amazon datasets. All hyperparameters were finalized based on validation performance, and results were reported using the test set. The experiments were conducted on one A6000 GPU.
@@ VI Experiment
  • RQ2: If the answer to RQ1 is ”Yes”, does IISAN-Versa exhibit a scaling effect with the integration of progressively larger LLM?
@@
VI-A IISAN-Versa Performance with larger text encoder (RQ1)
To validate RQ1, which is whether incorporating a larger text encoder can enhance the performance of IISAN-Versa, this section evaluates the performance of an asymmetrical IISAN-Versa (IISAN-VA) using the BERT-large as text encoder compared to a symmetrical IISAN-Versa (IISAN-VS) with a BERT-base encoder. The specifics of the models are detailed in Table II. In this section, we fix the image encoder as a ViT-base for a fair comparison, more image encoders will be further discussed in Section VI-B. We also use FFT and popular EPEFT approaches (Adapter, LoRA, and BitFit) as baselines. FFT and EPEFT methods were applied only to the symmetrical model due to the high computational cost of EPEFT and FFT. From Table VI-A, IISAN-VS demonstrates superior performance over traditional full fine-tuning and embedded PEFT approaches (Adapter, LoRA, and BitFit), as evidenced by its effectiveness across the three datasets.
The asymmetrical IISAN-Versa (IISAN-VA) further enhances performance with the larger text encoder. It achieves a gain of +6.31% in HR@10 and +6.12% in NDCG@10 for the Scientific dataset, +4.83% and +2.23% for the Instrument dataset, and +3.13% and +3.72% for the Office dataset. These improvements underscore IISAN-VA’s ability to effectively adapt to larger text encoder, leading to increased recommendation performance. This also demonstrates that compared to a BERT-base model, asymmetrical IISAN-Versa with a larger text encoder can generate a more powerful text representation for sequential recommendation tasks.
TABLE III: Performance comparison for IISAN-Versa. Note that the numbers indicated in this table are the percentage values. IISAN-VS and IISAN-VA represent symmetrical and asymmetrical IISAN-Versa. “*” denotes that the improvements of the best IISAN-VA compared with IISAN-VS (the second best) are significant at the level of 0.05 with paired T-test. The best results are in bold.
Dataset Metric FFT Adapter[14] LoRA[15] BitFit[16] IISAN-VS IISAN-VA Relative Improvement
Scientific HR@10 6.62 6.61 6.62 6.37 6.83 7.29 +6.31%
NDCG@10 4.06 3.91 4.09 3.76 4.14 4.41 +6.12%
[…]
NDCG@10 6.76 6.83 6.64 6.71 7.01 7.17 +2.23%
[…]
NDCG@10 4.58 4.85 4.78 4.87 4.92 5.11 +3.72%
[…]
@@ VI-A IISAN-Versa Performance with larger text encoder (RQ1)
TABLE IV: IISAN-Versa Encoders Scaling
Layers – Embedding Dim Model structures HR@10 NDCG@10
12 L 768 H + 12 L 768 H BERT-base+ViT-base 0.0683 0.0414
24 L 1024 H + 12 L 768 H BERT-large+ViT-base 0.0729 0.0441
28 L 3072 H + 12 L 768 H Gemma-7b-it+ViT-base 0.0732 0.0446
28 L 3072 H + 12 L 768 H Gemma-7b+ViT-base 0.0773 0.0475
32 L 4096 H + 12 L 768 H Mistral-7b+ViT-base 0.0753 0.0465
32 L 4096 H + 12 L 768 H Llama2-7b+ViT-base 0.0789 0.0481
32 L 4096 H + 12 L 768 H Llama2-chat-7b+ViT-base 0.0749 0.0453
32 L 4096 H + 12 L 768 H Llama3-8b+ViT-base 0.0778 0.0463
32 L 4096 H + 12 L 768 H Llama3-8b-it +ViT-base 0.0766 0.0474
(32 L 4096 H + 12 L 768 H) $\times$ 8 Mixtral-8x7b-it + ViT-base 0.0798 0.0480
(32 L 4096 H + 12 L 768 H) $\times$ 8 Mixtral-8x7b + ViT-base 0.0781 0.0473
80 L 8192 H + 12 L 768 H Llama3-70B-GPTQ-it +ViT-base 0.0801 0.0478
80 L 8192 H + 12 L 768 H Llama3-70B-GPTQ+ViT-base 0.0814 0.0494
12 L 768 H + 24 L 1024 H BERT-base+ViT-large 0.0688 0.0425
12 L 768 H + 32 L 1280 H BERT-base+ViT-huge 0.0682 0.0415
80 L 8192 H + 24 L 1024 H Llama3-70B-GPTQ+ViT-large 0.0756 0.0455
80 L 8192 H + 32 L 1280 H Llama3-70B-GPTQ+ViT-huge 0.0776 0.0464
@@ VI-B Scaling Effect (RQ2)
As depicted in Table IV, IISAN-Versa equipped with the largest scale of Llama3-70b-GPTQ model demonstrated the most superior performance. The trend of these results is visualized in Figure 4, which shows the overall trajectory is generally upward. Moreover, the average performance within each group significantly improves with each scale increase, supporting the hypothesis that IISAN-Versa follows a scaling effect for text encoders. This observation validates the IISAN-Versa framework’s capability to leverage advancements from pre-trained foundation models in the NLP domain. We can observe that, with the exception of the Mixtral 8x7b model, most LLMs with instruction tuning typically do not outperform those without it. This reflects that the LLM without instruction tuning may generate a more powerful representation.
TABLE V: Performance comparison of IISAN-Versa on a multimodal public benchmark Microlens. In this experiment, IISAN-VS and IISAN-VA adopt BERT-base and BERT-large respectively. “*” denotes that the improvements of the best models compared with previous state-of-the-art GRU4Recv are significant at the level of 0.05 with paired T-test.
@@
On the other hand, as shown in Table IV, while integrating the ViT-large model for the BERT-base model marginally enhances the IISAN-Versa’s performance by less than 1%, upgrading to the ViT-huge model does not yield any further improvements. For the Llama3-70b-GPTQ, further enlarging the vision encoder, the performance slightly dropped. This contrasts with the text encoders, where similar scaling scales like the BERT-large have led to performance gains exceeding 6%. This disparity might stem from two potential factors: (1) the text encoder’s predominance as the primary modality means that enhancements to the image encoder do not substantially elevate overall performance, and (2) the implemented image encoders may not be sufficiently large and effective to significantly better the representation compared to the baseline ViT-base model. The first hypothesis is also supported by an ablation study highlighting the text modality’s greater significance, whereas the second hypothesis remains speculative in the absence of access to larger vision transformers like the ViT-22B.
(Answer to RQ2) IISAN-Versa complies with the scaling effect of text encoders, confirming its potential to inherit the benefits of advanced pre-trained text transformer models. This also indicates that researchers can explore LLMs not only as powerful text generators but also as effective text encoders in future works, thus reaching more comprehensive representations for recommender systems. Further analysis suggests that substantial gains in overall performance are unlikely from merely increasing the size of the vision transformer. Therefore, based on current observations, it is advisable to continue utilizing the base or large versions of the vision transformer while prioritizing the expansion of the text encoder.

### v2

@@ III-E LayerDrop
Additionally, to further enhance network efficiency and address issues of layer redundancy, we introduce LayerDrop techniques, similar to those in [36, 14], to reduce the number of SAN blocks and align asymmetrical MFM. Specifically, for symmetrical MFM, we group two transformer blocks together and drop the first hidden state to the SANs, which can save half the number of SANBs. In the case of asymmetrical MFM, to address the imbalance between the layers of LLM and vision encoders, we apply a simple yet effective group LayerDrop technique that groups $k$ layers evenly from the hidden states of the LLM. Suppose there are $L^{\text{t}}$ and $L^{\text{v}}$ layers in the text encoder and image encoder, respectively. In scenarios of asymmetrical adaptation, the text encoder is an LLM, thus $L^{\text{t}}\gg L^{\text{v}}$. For image encoders, we continue using the same LayerDrop technique as with symmetrical structures, where we drop half the number of layers. In this paper, we design to select the largest $k$ to ensure the maximum range of cross-layer interaction within the entire LLM. The number $k$ can be represented as follows:
$\max_{k}\left(L^{\text{t}}-k\cdot\frac{L^{\text{v}}}{2}\geq 1\right)$ (5)
We compute $\mathrm{step}=\bigl{\lceil}\tfrac{L}{L^{v}}\bigr{\rceil}$ and select layers as shown in Algorithm 1. The specific selections for various $L$ (with $L^{v}=6$) are presented in Table I. In Section VI-E, we further investigate various LayerDrop schemes, showcasing the performance impact of dropping different numbers of blocks in both symmetrical and asymmetrical structures.
@@ III-G Training Objectives
In terms of training details, we exploit the in-batch debiased Cross-Entropy loss function $\mathcal{L}_{CE}$ [2, 18] widely adopted in both academic literature and industrial systems.
@@ IV Efficiency Analysis
To simplify the theoretical analysis, we consider a scenario involving large foundation models comprising $n$ transformer blocks, where the $i$-th transformer block is denoted by $TRM_{i}$. For each transformer block, all PEFT methods, including Adapter, LoRA, and IISAN, will place a trainable PEFT module denoted by peftpeft. We assume that $TRM\gg peft$ according to the definition of parameter efficiency.
@@ V Experiment Setup
TABLE II: Dataset Description
@@ V-A Dataset
In this study, each item in a user sequence is represented by its raw modality information. To assess our methods on items that include both image and text modalities, we utilize the Amazon review datasets [45]. Specifically, we employ three widely used datasets from this collection: Industrial and Scientific,” Musical Instruments,” and Office Products.” These datasets were adopted following the preprocessing steps outlined in prior works [14, 2, 10, 46]. Furthermore, to explore the performance of IISAN-Versa compared with other recommendation approaches with a fair comparison, we also incorporate MicroLens [18], a public multimodal video recommendation benchmark that includes raw text, images, and video content. Since MicroLens is a dataset rich in modality information, we also explore multimodal text scenarios within this public benchmark. To improve the evaluation scale, we additionally sampled one million interactions from the publicly available NINERec dataset [47].
@@ V-B Multimodal Text Generation
To explore the potential of IISAN-Versa in adapting multiple scenarios, such as multimodal text scenarios containing cover text and corresponding video texts, we utilize the pretrained LLaVA model to generate captions for image and video content from the public benchmark MicroLens datasets, as illustrated in Figure 3. For this process, we input the cover image and extract three evenly spaced frames from each video. In terms of caption models, we use llava-hf/llava-v1.6-mistral-7b-hf” for cover image captioning. For video captioning, we apply liuhaotian/llava-v1.6-34b” due to the higher complexity of videos; a larger caption model provides a more accurate representation of the video content. Our video captioning strategies follow the preprocessing methods used in Video generation OpenSora [^3] https://github.com/hpcaitech/Open-Sora. To enhance LLaVA’s ability to produce effective representations for postprocessing, we incorporate in-context learning by including concrete examples in the prompt. We adopt the prompt from the Sora project [^4] https://openai.com/sora for video prompts and use the prompt from DALLE2555https://openai.com/index/dall-e-2 for cover image prompts.
@@
Based on previous research [48, 2, 10, 14], our methodology employs a leave-one-out evaluation strategy. In this approach, the final item in the interaction sequence is reserved for testing, the penultimate item is used for validation, and all preceding items are utilized for training. We evaluate our model using HR@10 (Hit Ratio) and NDCG@10 (Normalised Discounted Cumulative Gain), consistent with the evaluation methodologies in [48, 2, 10], as the key performance metrics. Unless stated otherwise, all results reported refer to the test set. Furthermore, it is crucial to mention that the predicted item is evaluated against the complete set of items.
V-D Implementation Details.
The pretrained models from the Huggingface platform [^6] https://huggingface.co/, detailed in Table III, are employed as text and image encoders in this research. Note that due to the limited computing resources, we adopt the Llama-3-70b with GPTQ 4-bit quantization. We experiment with hidden representation dimensions for the sequential encoder from {32, 64, 128} and set it to 64. We fixed the sequential encoder to include 2 Transformer blocks and 2 attention heads following [14, 2].
TABLE III: Description of the pre-trained models used in this research. L denotes the number of Transformer blocks, H represents the hidden dimension of the transformer layer, B indicates the number of parameters in billions, and M indicates the number of parameters in millions.
@@
Text BERT-tiny L=2, H=128 4.4M
BERT-base L=12, H=768 110M
@@
EVA-ViT L=48, H=5120 11B
Using the Adam optimizer without weight decay, we tested learning rates from 1e-5 to 1e-3 while maintaining a dropout probability of 0.1. We explored batch sizes from {8, 16, 32, 64, 512}, selecting the largest viable option to optimize GPU memory, and searched for SANB hidden dimensions from {8, 16, 32, 64, 128, 512, 1024, 2048}. Hyperparameters were searched separately for the MicroLens and Amazon datasets. All hyperparameters were finalized based on validation performance, and results were reported using the test set. The experiments were conducted on one A6000 GPU.
@@ VI Experiment
  • RQ2: If the answer to RQ1 is ”Yes”, does IISAN-Versa exhibit a scaling effect with the integration of progressively larger LLMs?
@@
  • RQ7: How does the IISAN-Versa perform with different hyper-parameters?
VI-A IISAN-Versa Performance with larger text encoders (RQ1)
TABLE IV: Performance (%) comparison across methods on three datasets. Best results in bold, second-best underlined. “*” denotes significant improvement over the second-best (IISAN-VS) at 0.05 level with paired T-test.
Method Scientific Instrument Office
H@10 N@10 H@10 N@10 H@10 N@10
FFT 6.62 4.06 8.76 6.76 6.30 4.58
Adapter 6.61 3.91 8.82 6.83 6.65 4.85
LoRA 6.62 4.09 8.43 6.64 6.55 4.78
BitFit 6.37 3.76 8.65 6.71 6.78 4.87
IISAN-VS 6.83 4.14 9.06 7.01 6.80 4.92
IISAN-VA 7.29 4.41 9.52 7.17 7.02 5.11
To validate RQ1, which is whether incorporating a larger text encoder can enhance the performance of IISAN-Versa, this section evaluates the performance of an asymmetrical IISAN-Versa (IISAN-VA) using the BERT-large as text encoder compared to a symmetrical IISAN-Versa (IISAN-VS) with a BERT-base encoder. The specifics of the models are detailed in Table III. In this section, we fix the image encoder as a ViT-base for a fair comparison, more image encoders will be further discussed in Section VI-B. We also use FFT and popular EPEFT approaches (Adapter, LoRA, and BitFit) as baselines. FFT and EPEFT methods were applied only to the symmetrical model due to the high computational cost of EPEFT and FFT. From Table IV, IISAN-VS demonstrates superior performance over traditional full fine-tuning and embedded PEFT approaches (Adapter, LoRA, and BitFit), as evidenced by its effectiveness across the three Amazon datasets.
The asymmetrical IISAN-Versa (IISAN-VA) further enhances performance with the larger text encoder. It achieves a gain of +6.31% in HR@10 and +6.12% in NDCG@10 for the Scientific dataset, +4.83% and +2.23% for the Instrument dataset, and +3.13% and +3.72% for the Office dataset. These improvements underscore IISAN-VA’s ability to effectively adapt to larger text encoders, leading to increased recommendation performance. This also demonstrates that compared to a BERT-base model, asymmetrical IISAN-Versa with a larger text encoder can generate a more powerful text representation for sequential recommendation tasks.
@@ VI-A IISAN-Versa Performance with larger text encoders (RQ1)
TABLE V: IISAN‑Versa Encoders Scaling.
Model structure HR@10 NDCG@10
BERT‑base + ViT‑base 0.0683 0.0414
BERT‑large + ViT‑base 0.0729 0.0441
Gemma‑7b‑it + ViT‑base 0.0732 0.0446
Gemma‑7b + ViT‑base 0.0773 0.0475
Mistral‑7b + ViT‑base 0.0753 0.0465
Llama2‑7b + ViT‑base 0.0789 0.0481
Llama2‑chat‑7b + ViT‑base 0.0749 0.0453
Llama3‑8b + ViT‑base 0.0778 0.0463
Llama3‑8b‑it + ViT‑base 0.0766 0.0474
Mixtral‑8×7b‑it + ViT‑base 0.0798 0.0480
Mixtral‑8×7b + ViT‑base 0.0781 0.0473
Llama3‑70B‑GPTQ‑it + ViT‑base 0.0801 0.0478
Llama3‑70B‑GPTQ + ViT‑base 0.0814 0.0494
BERT‑base + ViT‑large 0.0688 0.0425
BERT‑base + ViT‑huge 0.0682 0.0415
BERT‑base + EVA‑ViT 0.0746 0.0457
Llama3‑70B‑GPTQ + ViT‑large 0.0756 0.0455
Llama3‑70B‑GPTQ + ViT‑huge 0.0776 0.0464
Llama3‑70B‑GPTQ + EVA‑ViT 0.0762 0.0469
Llama3‑70B‑GPTQ + ViT‑tiny 0.0780 0.0478
BERT‑tiny + EVA‑ViT 0.0744 0.0449
TABLE VI: Performance comparison of IISAN-Versa on a large dataset over 1M interactions. “*” denotes that the improvements are significant at the level of 0.05 with paired T-test.
Class Model HR@10 NDCG@10
MM VideoRec Full finetuning 0.0098 0.0048
Adapter 0.0097 0.0048
IISAN-Versa IISAN-VS (ours) 0.0105 0.0052
IISAN-VA (ours) 0.0111 0.0054
@@ VI-B Scaling Effect (RQ2)
As depicted in Table V, IISAN-Versa equipped with the largest scale of Llama3-70b-GPTQ model demonstrated the most superior performance. The trend of these results is visualized in Figure 4, which shows the overall trajectory is generally upward. Moreover, the average performance within each group significantly improves with each scale increase, supporting the hypothesis that IISAN-Versa follows a scaling effect for text encoders. This observation validates the IISAN-Versa framework’s capability to leverage advancements from pre-trained foundation models in the NLP domain. We can observe that, with the exception of the Mixtral 8x7b model, most LLMs with instruction tuning typically do not outperform those without it. This reflects that the LLM without instruction tuning may generate a more powerful representation.
TABLE VII: Performance comparison of IISAN-Versa on a multimodal public benchmark MicroLens. In this experiment, IISAN-VS and IISAN-VA adopt BERT-base and BERT-large respectively. “*” denotes that the improvements of the best models compared with previous state-of-the-art GRU4Recv are significant at the level of 0.05 with paired T-test.
@@
As shown in Table V, replacing ViT-base with ViT-large or ViT-huge yields only minor improvements, and in some cases even reduces performance—especially when using strong text encoders like Llama3-70B-GPTQ. To explore further, we include EVA-CLIP-18B which contains one of the largest publicly available vision transformers. While it significantly boosts performance when paired with lightweight text models, its impact is limited when the text encoder is already strong. This suggests that once strong textual representations are in place, the value of visual features diminishes—a trend aligned with prior findings [10, 2] and supported by our ablation studies showing the dominant role of text in multimodal recommendation. To further strengthen our analysis, we include extreme architectural asymmetry—where the text encoder is significantly larger or smaller than the vision encoder—demonstrating the robustness of our IISAN-Versa.
Beyond model-wise scaling, we evaluate IISAN-Versa on the large-scale NINERec dataset (over 1M interactions) as shown in Table VI. IISAN-VS and IISAN-VA achieve competitive performance over FFT and Adapter, highlighting the practical effectiveness of our approach.
(Answer to RQ2) IISAN-Versa complies with the scaling effect of text encoders, confirming its potential to inherit the benefits of advanced pre-trained text transformer models. This also indicates that researchers can explore LLMs not only as powerful text generators but also as effective text encoders in future works, thus reaching more comprehensive representations for recommender systems. Further analysis shows that a large vision encoder, such as EVA-ViT, can effectively compensate when the text encoder is relatively weak, leading to noticeable performance gains. However, its impact diminishes once the text encoder is already strong, suggesting that visual enhancements offer limited benefit in such cases. Therefore, scaling the vision encoder is most beneficial in asymmetric setups with lightweight text models, but less effective when paired with high-capacity text encoders.
[Uncaptioned image] Figure 5: Comparison of GPU Memory Usage and Performance Between SASRec and IISAN-Versa.
TABLE VIII: Efficiency comparison with optimal SASRec and IISAN-Versa. GM represents GPU memory (GB), TT represents Training time (seconds) per epoch. Model GM TT IISAN-Versa 4.38 51 SASRec 4.96 9